A college freshman borrowed P2,000 from a bank for his tuition fee and promised to pay the amount for one year. He received only the amount of P1,920 after the bank collected the advance interest of P80.00. What was the rate of discount?


1
Expert's answer
2022-03-14T09:56:17-0400

let the amount borrowed (P) = 2000

Let the amount received (A) = 1920

Let the advance interest = 80

"80=P*R*T"

80"=2000*R*1"

"R=80\/2000= 0.04"

The percentage of the rate of discount = "0.04*100= 4.00%"

The rate of discount is 4.00%